### v1

# Noah Porcelli

Abstract
We prove that (under appropriate orientation assumptions), the action of a Hamiltonian homeomorphism $\phi$ on the cohomology of a relatively exact Lagrangian fixed by $\phi$ is the identity. This extends results of Hu-Lalonde-Leclercq [7] and the author [10] in the setting of Hamiltonian diffeomorphisms. We also prove a similar result regarding the action of $\phi$ on relative cohomology.
1 Introduction
Let $(M,\omega)$ be a symplectic manifold, and $L\subseteq M$ a Lagrangian submanifold. In [13], Yau studied the following question:
Question 1.1.
What diffeomorphisms $\tilde{\phi}$ of $L$ can be extended to a Hamiltonian diffeomorphism $\phi$ on $M$?
This question has been studied in various places. For example, some partial classification results in the case $L$ is a monotone 2-torus have been obtained in [13, 1]; in particular, even in simple cases $\tilde{\phi}$ can act non-trivially on $H^{*}(L)$. A more comprehensive history of this question can be found in the introduction in [1].
We now assume (for the rest of the paper) that $M$ is compact or Liouville, and that $L\subseteq M$ is a compact, relatively exact Lagrangian. We also assume all Hamiltonians are compactly supported: this is without any loss of generality, as can be seen by using an appropriate cut-off function supported away from the image of $L$ under the flow of this Hamiltonian.
In this setting, strong constraints have been obtained for $\tilde{\phi}$ to extend:
Theorem 1.2.
Let $\tilde{\phi}\mathrel{\mathop{:}}L\to L$ be a diffeomorphism, and assume that there is a Hamiltonian diffeomorphism $\phi\mathrel{\mathop{:}}M\to M$ extending $\tilde{\phi}$.
  1. 1.
    ([7]): $\phi$ acts as the identity on $H^{*}(L;\mathbb{Z}/2)$.
  2. 2.
    ([7, 10]): If $L$ is Spin, $\phi$ acts as the identity on $H^{*}(L;\mathbb{Z})$.
  3. 3.
    ([12]): If $L$ is oriented, $\phi$ is orientation-preserving.
Further results regarding the action of $\phi$ on string topology as well as some generalised cohomology theories were obtained in [10]. Motivated by these, one can conjecture:
Conjecture 1.3.
Consider the space $\mathcal{G}$ of Hamiltonian diffeomorphisms $\phi\mathrel{\mathop{:}}M\to M$ sending $L$ to itself. Then $\mathcal{G}$ is contractible.
The case where $L$ is the zero section in $M=T^{*}L$ would follow from the strong nearby Lagrangian conjecture.
1.1 Extension to the $C^{0}$ setting
Definition 1.4.
A Hamiltonian homeomorphism is a homeomorphism $\theta\mathrel{\mathop{:}}M\to M$ which is a $C^{0}$ limit of Hamiltonian diffeomorphisms.
It is known due to work of Eliashberg-Gromov [3, 4] that any Hamiltonian homeomorphism which is also a diffeomorphism, is in fact a Hamiltonian diffeomorphism. In general however the geometry of Hamiltonian homeomorphisms is much less understood than those of Hamiltonian diffeomorphisms.
The dynamical behaviour of Hamiltonian homeomorphisms is very different to Hamiltonian diffeomorphisms: for example, the Arnold conjecture fails (badly) for Hamiltonian homeomorphisms [2].
While Question 1.1 has not been studied in the $C^{0}$ setting, a similar question has been studied very recently in [8].
Here we extend some of the results of Theorem 1.2 to the $C^{0}$ setting (it would be interesting to extend the other results to the $C^{0}$ setting too).
Theorem 1.5.
Let $\tilde{\phi}\mathrel{\mathop{:}}L\to L$ be a homeomorphism, and assume that there is a Hamiltonian homeomorphism $\phi\mathrel{\mathop{:}}M\to M$ extending $\tilde{\phi}$.
  1. 1.
    $\phi$ acts as the identity on $H^{*}(L;\mathbb{Z}/2)$.
  2. 2.
    If $L$ is Spin, $\phi$ acts as the identity on both cohomology $H^{*}(L;\mathbb{Z})$ and complex $K$-theory $K^{*}(L)$.
Theorem 1.5 holds for various other generalised cohomology theories $R^{*}$ (such as real $K$-theory $KO^{*}$) provided the condition [10, Proposition 1.14] holds for $R^{*}$. Examples of diffeomorphisms of closed manifolds which act as the identity but not on (real or complex) $K$-theory are given in [10, Appendices A & B].
The proof of Theorem 1.5 approximates $\phi$ with a $C^{0}$-close Hamiltonian diffeomorphism, and proceeds by adapting the strategy of [10] to this approximation.
Remark 1.6.
If $M$ is Liouville and $L$ is exact, it would follow from the nearby Lagrangian conjecture that if $\tilde{\phi}$ is as in Theorem 1.5, then $\tilde{\phi}$ is homotopic to a diffeomorphism $L\to L$ which extends to a Hamiltonian diffeomorphism of $M$; Theorem 1.5 would then follow from Theorem 1.2.
1.2 Extension to the relative setting
Assume $\tilde{\phi}\mathrel{\mathop{:}}L\to L$ is a homeomorphism which extends to a Hamiltonian homeomorphism $\phi\mathrel{\mathop{:}}M\to M$. $\phi$ induces a map of pairs $(M,L)\to(M,L)$, and hence a map on relative cohomology. Even though $\phi$ is homotopic to the identity on $M$ (just as a map, not necessarily of pairs), and $\tilde{\phi}$ acts as the identity on cohomology, it does not immediately imply the map induced on relative cohomology is the identity, though Conjecture 1.3 would imply this. Nevertheless:
Theorem 1.7.
Let $\tilde{\phi}\mathrel{\mathop{:}}L\to L$ be a homeomorphism, and assume that there is a Hamiltonian homeomorphism $\phi\mathrel{\mathop{:}}M\to M$ extending $\tilde{\phi}$.
  1. 1.
    $\phi$ acts as the identity on $H^{*}(M,L;\mathbb{Z}/2)$.
  2. 2.
    If $L$ is Spin, $\phi$ acts as the identity on both $H^{*}(M,L;\mathbb{Z})$ and $K^{*}(M,L)$.
The proof is similar to the proof of Theorem 1.5, but whereas the proof of Theorem 1.5 uses a moduli space of holomorphic curves with domain $Z_{C}$, $\mathcal{P}_{C}$, along with the evaluation map on the boundary of $Z_{C}$, $ev\mathrel{\mathop{:}}\mathcal{P}\times\partial Z_{C}\to M$, the proof of Theorem 1.7 uses the evaluation map on the whole of $Z_{C}$, $ev\mathrel{\mathop{:}}\mathcal{P}\times Z_{C}\to M$.
1.3 Acknowledgements
The author thanks Amanda Hirschi, Ivan Smith and Yuhan Sun for helpful conversations, and Jack Smith for helpful conversation and useful comments on an earlier draft. The author is supported by the Engineering and Physical Sciences Research Council [EP/W015889/1].
2 Smooth approximation
We first choose a complete Riemannian metric on $M$, and let $03\varepsilon>$ 0 be its injectivity radius. Write $d$ for the Riemannian distance function.
Now choose a smooth Hamiltonian diffeomorphism $\psi^{1}\mathrel{\mathop{:}}M\to M$ which is sufficiently $C^{0}$-close to $\phi$ that for all $x$ in $M$, $d(\psi^{1}(x),\phi(x))\leq\varepsilon$. We extend $\psi^{1}$ to a 1-parameter family $\{\psi^{t}\}_{t\in[0,1]}$ of Hamiltonian diffeomorphisms such that $\psi^{0}$ is the identity. We write $L^{\prime}$ for $\psi^{1}(L)$.
Let $U$ be a tubular neighbourhood of $L$ and $\pi\mathrel{\mathop{:}}U\to L$ the projection map. We choose $U$ small enough that for all $x$ in $U$,
$d(x,\pi(x))\leq\varepsilon$ (1)
We may choose $\psi^{1}$ $C^{0}$-close enough to $\phi$ such that $L^{\prime}$ lives in $U$, and generically so that $L$ and $L^{\prime}$ intersect transversally.
Lemma 2.1.
The maps $\phi,\pi\circ\psi^{1}\mathrel{\mathop{:}}L\to L$ are homotopic.
Proof.
For $x$ in $L$, since $\pi\circ\phi(x)=\phi(x)$, by the triangle inequality $d(\phi(x),\pi\circ\psi^{1}(x))\leq 3\varepsilon$. Thus there is a unique small geodesic between $\phi(x)$ and $\pi\circ\psi^{1}(x)$; moving along this geodesic for each $x$ gives the required homotopy. ∎
3 Moduli spaces
In this section, we construct moduli spaces $\mathcal{M}$ and $\mathcal{P}_{r}$ for $r\geq 0$, similarly to [10, Sections 4.1 & 5.3] and [6].
Let $D$ be a smooth convex subdomain of $\mathbb{C}$, lying inside $\mathbb{R}+i[0,1]$ and containing $[-1,1]+i[0,1]$.
Definition 3.1.
We define $Z_{-}\mathrel{\mathop{:}}=D\cup\left([0,\infty)+i[0,1]\right)$.
For $r\geq 0$, we define $Z_{r}$ to be the union
$Z_{r}\mathrel{\mathop{:}}=\left(D-r\right)\cup\left([-r,r]+i[0,1]\right)\cup\left(D+r\right)$
See below for pictures:
{tikzpicture}
$Z_{-}$ and $Z_{r}$ are both smooth convex subdomains of $\mathbb{C}$.
Recall the (topological) energy of a $J$-holomorphic map $u\mathrel{\mathop{:}}\Sigma\to M$ from a Riemann surface $\Sigma$ is the quantity
$E(u)\mathrel{\mathop{:}}=\int_{\Sigma}u^{*}\omega$
Definition 3.2.
For $r\geq 0$, we define $\mathcal{P}_{r}$ to be the space of $J$-holomorphic maps $u\mathrel{\mathop{:}}Z_{r}\to M$ with finite energy, with moving Lagrangian boundary conditions:
$u(z)\in\psi^{\operatorname{Im}(z)}(L)\,\textrm{ for all }\,z\in\partial Z_{r}$ (2)
We define $\mathcal{M}$ to be the space of $J$-holomorphic maps $u\mathrel{\mathop{:}}Z_{-}\to M$ which have finite energy, satisfying (2) (with $Z_{r}$ replaced by $Z_{-}$).
In particular, if $x\in\mathbb{R}$, each of these moduli spaces sends $x$ to $L$ and $x+i$ to $L^{\prime}$ (assuming $x,x+i$ lie in the relevant domain).
Note that we do not quotient by any automorphisms in the definitions of these moduli spaces; this does not interfere with compactness, by the argument in [10, Lemma 4.5].
[10, Lemma 4.3] shows there is an a priori uniform upper energy bound on the energy of any $u\in\mathcal{M}$ or $u\in\mathcal{P}_{r}$, independent of $r$ (dependent however on the choice of $\{\psi^{t}\}_{t}$).
Using relative exactness, Gromov compactness [9, Theorem 4.6.1] implies that
Proposition 3.3.
Each $\mathcal{P}_{r}$, as well as $\mathcal{M}$, is compact.
Furthermore for any sequence $u_{n}\in\mathcal{P}_{r_{n}}$ where $r_{n}\to\infty$, after restricting to some subsequence, there is some $u\in\mathcal{M}$ such that for any compact subset $K\subseteq Z_{-}$,
$u_{n}(\cdot-C_{n})|_{K}\to u|_{K}$
converges uniformly.
Though we could generically choose $J$ so that these moduli spaces are transversally cut out, this would be unnecessary for our purposes.
Proposition 3.4.
Let $C\geq 0$ and $x\in[-C,C]$, and let $ev_{x}\mathrel{\mathop{:}}\mathcal{P}_{C}\to L$ send $u$ to $u(x)$. Then under the orientation conditions of Theorem 1.5 for $R^{*}=H^{*}(\cdot;\mathbb{Z}/2),H^{*}(\cdot;\mathbb{Z})$ or $K^{*}$, the induced map $ev^{*}_{x}\mathrel{\mathop{:}}R^{*}(L)\to R^{*}(\mathcal{P}_{C})$ is injective.
This is proved in [6, Theorem 3.1] in the case $R^{*}$ is (a Čech model for) $H^{*}(\cdot;\mathbb{Z}/2)$ and [5, Theorem 2.12] in the other cases. Since $\mathcal{P}_{C}$ can in generally be extremely singular (and in particular, not necessarily homotopy equivalent to a CW complex), we must be careful with exactly what model of $R^{*}(\mathcal{P}_{C})$ we use; see [5, Section 2.1] for a discussion about this point. However all reasonable models for $R^{*}(\cdot)$ agree when applied to $L$, since $L$ admits the structure of a finite CW complex, so this does not affect how we apply Proposition 3.4.
4 Pinching
For $r\geq 0$, $u\in\mathcal{P}_{r}$ and $x\in[-r,r]$, we define as shorthand
$d_{u}(x)\mathrel{\mathop{:}}=\sup_{t\in[0,1]}d\left(u(x),u(x+ti)\right)$
Proposition 4.1.
There is some $C\geq 0$ and $x_{0}\in[-C,C]$ such that for all $u$ in $\mathcal{P}_{C}$,
$d_{u}(x_{0})\leq\varepsilon$
This should be viewed as a version of Gromov compactness; indeed we will deduce Proposition 4.1 from Proposition 3.3. First we need a lemma:
Lemma 4.2.
There is some $\lambda>0$ such that for all $v$ in $\mathcal{M}$,
$d_{v}(\lambda)\leq\frac{\varepsilon}{2}$
Proof.
Since we assumed $L$ and $L^{\prime}$ intersect transversally, for fixed $v\in\mathcal{M}$ this holds for $\lambda$ sufficiently large, by the standard exponentian decay estimate for holomorphic curves on their strip-like ends; see e.g. [11]; we may choose $\lambda>0$ large enough that this holds for all $v\in\mathcal{M}$ since $\mathcal{M}$ is compact. ∎
Proof of Proposition 4.1.
Let $\lambda$ be as in Lemma 4.2. We show that Proposition 4.1 holds for $C$ sufficiently large and $x_{0}=\lambda-C$.
Suppose not. There is then a sequence $C_{n}\to\infty$ and $u_{n}\in\mathcal{P}_{C_{n}}$, such that for all $n$,
$d_{u_{n}}(\lambda-C_{n})>\varepsilon$ (3)
Applying Proposition 3.3 to $K=\lambda+i[0,1]\subseteq Z_{-}$, we find some $u\in\mathcal{M}$ such that
$u_{n}(\cdot-C_{n})|_{K}\to u|_{K}$ (4)
converges uniformly. Therefore
$d_{u_{n}}(\lambda-C_{n})\to d_{u}(\lambda)$ (5)
The right hand side of (5) is $\leq\varepsilon/2$ by Lemma 4.2, whereas the right hand side is $>\varepsilon$ for all $n$, by (3), giving a contradiction. ∎
5 Proof of Theorem 1.5
Let $C$ and $x_{0}$ be as in Proposition 4.1 and let $ev_{x_{0}},ev_{x_{0}+i}\mathrel{\mathop{:}}\mathcal{P}_{C}\to L$ be the maps which evaluate at $x_{0}$ and $x_{0}+i$ respectively
Lemma 5.1.
The following diagram commutes up to homotopy:
[…]
Proof.
We write down an explicit homotopy; this can be thought of as “evaluating around the boundary of $Z_{C}$ from $x_{0}$ to $x_{0}+i$”. Let $\gamma\mathrel{\mathop{:}}[0,1]\to\partial Z_{C}$ be any path from $x_{0}$ to $x_{0}+i$, and let $r_{t}=\operatorname{Im}\gamma(t)$ for $t\in[0,1]$. Then we define a homotopy $H\mathrel{\mathop{:}}[0,1]\times\mathcal{P}_{C}\to L^{\prime}$ to send $(t,u)$ to
$\psi^{1}\circ\left(\psi^{r_{t}}\right)^{-1}\circ u\circ\gamma(t)$
Then $H(0,\cdot)=\psi^{1}\circ ev_{x_{0}}$ and $H(1,\cdot)=ev_{x_{0}+i}$. ∎
Using Proposition 4.1, we may replace $\phi$ with $\psi^{1}$:
Corollary 5.2.
The following diagram commutes up to homotopy:
[…]
Proof.
By Proposition 4.1 and (1), for all $u\in\mathcal{P}_{C}$, $\pi\circ ev_{x_{0}+i}(u)$ and $ev_{x_{0}}(u)$ are of distance at most $2\varepsilon$ from each other. Since $2\varepsilon$ is less than the injectivity radius, the following diagram commutes, via a homotopy which follows the unique small geodesic between the two endpoints:
[…]
Combined with (6) and Lemma 2.1, this implies the result. ∎
Proof of Theorem 1.5.
Let $R^{*}$ be the relevant cohomology theory. Then by applying $R^{*}$ to (7) and using Proposition 3.4, we see that for any $\alpha\in R^{*}(L)$, $\phi^{*}(\alpha)=\alpha$. ∎
6 Proof of Theorem 1.7
Let $R$ be the cohomology theory $H^{*}(\cdot;\mathbb{Z}/2),H^{*}(\cdot;\mathbb{Z})$ or $K^{*}$, and assume the relevant orientation condition from Theorem 1.7 holds.
We let $C$ and $x_{0}$ be as in Proposition 4.1. We define a family of maps $\{\phi^{t}\}_{t\in[0,1]}$ to be given by $\phi^{t}=\psi^{2t}$ for $0\leq t\leq\frac{1}{2}$ and to follow the unique small geodesic between $\psi^{1}(x)$ and $\phi(x)$ for $x\in M$ and $\frac{1}{2}\leq t\leq 1$. Then $\phi^{1}=\phi$ and $\phi^{0}$ is the identity.
Consider the sweep-out map $S\mathrel{\mathop{:}}L\times[0,1]\to M$, sending $(x,t)$ to $\phi^{t}(x)$. This sends $L\times\{0,1\}$ to $L$ and therefore a map of pairs $(L\times[0,1],L\times\{0,1\})\to(M,L)$; applying $R^{*}$ and applying the suspension isomorphism gives a map
$\sigma\mathrel{\mathop{:}}R^{*+1}(M,L)\to R^{*}(L)$
Let $\partial\mathrel{\mathop{:}}R^{*}(L)\to R^{*+1}(M,L)$ be the boundary map.
Lemma 6.1.
The two maps
$\phi^{*},(Id+\partial\circ\sigma)\mathrel{\mathop{:}}R^{*}(M,L)\to R^{*}(M,L)$
are equal.
Proof.
Let $\overline{M}=M\cup_{L\times\{0\}}L\times[0,1]$ be the mapping cylinder of the inclusion $L\hookrightarrow M$. Then $\phi$ induces a map of pairs $(\overline{M},L\times\{1\})$ to itself, and the action of this map on $R^{*}$ is the same as that of the action of $\phi$ on $(M,L)$.
We define another map of pairs $\theta$ from $(\overline{M},L\times\{1\})$ to itself: $\theta$ sends $x\in M$ to $x$ and $(x,t)\in L\times[0,1]$ to $(\phi^{t}(x),t)$. By construction, the map induced by $\theta$ on $R^{*}(M,L)$ is $Id+\partial\circ\sigma$, so it remains to show that $\theta$ and $\phi$ are homotopic as maps of pairs.
Define $H\mathrel{\mathop{:}}\overline{M}\times[0,1]_{s}\to\overline{M}$ to send $(x,s)\in M\times[0,1]$ to $\phi^{s}(x)$ and $(x,t,s)\in L\times[0,1]^{2}$ to
$\left(\phi^{\min\{s+t,1\}}(x),t\right)$
Then $H(\cdot,0)=\theta$, $H(\cdot,1)=\phi$ and $H(\cdot,t)$ sends $L\times\{1\}$ to itself for all $t$; together these imply that $H$ is the desired homotopy. ∎
We a map define a map $T\mathrel{\mathop{:}}\mathcal{P}_{C}\times[0,1]\to M$ as follows. Let $\gamma\mathrel{\mathop{:}}[0,1]\to\partial Z_{C}$ be a path from $x_{0}$ to $x_{0}+i$ and let $r_{t}=\operatorname{Im}\gamma(t)$. Then we define $T(u,t)$ to be $u\circ\gamma(2t)$ for $0\leq t\leq\frac{1}{2}$ and to follow the unique small geodesic between $ev_{x_{0}+i}(u)$ and $\phi\circ(\psi^{1})^{-1}\circ ev_{x_{0}+i}(u)$ for $\frac{1}{2}\leq t\leq 1$.
We further define a map $F\mathrel{\mathop{:}}\mathcal{P}_{C}\times[0,1]\to L\times[0,1]$ by
$F(u,t)\mathrel{\mathop{:}}=\begin{cases}\left(\left(\psi^{r_{2t}}\right)^{-1}\circ u\circ\gamma(2t),t\right)&\textrm{ if }0\leq t\leq\frac{1}{2}\\ \left(\left(\psi^{1}\right)^{-1}\circ ev_{x_{0}+i}(u),t\right)&\textrm{ if }\frac{1}{2}\leq t\leq 1\end{cases}$
Lemma 6.2.
The following diagram commutes up to homotopy (relative to $\mathcal{P}_{C}\times\{0,1\}$):
[…]
Proof.
We define a map $S^{\prime}\mathrel{\mathop{:}}L\times[0,1]\to M$ to send $(x,t)$ to $\psi^{r_{2t}}(x)$ if $0\leq t\leq\frac{1}{2}$ and to follow the unique small geodesic between $\psi^{1}(x)$ and $\phi(x)$ for $\frac{1}{2}\leq t\leq 1$. By construction, the following diagram commutes:
[…]
$S^{\prime}$ and $S$ are homotopic (relative to $L\times\{0,1\}$) since they are the same up to a reparametrisation of $[0,\frac{1}{2}]$.
Lemma 6.3.
The sweep-out map $\sigma$ vanishes.
Proof.
Proposition 3.4 implies that the map
$R^{*}(L\times[0,1],L\times\{0,1\})\to R^{*}(\mathcal{P}_{C}\times[0,1],\mathcal{P}_{C}\times\{0,1\})$
is injective, so it suffices to show that $T$ is homotopic (through maps sending $\mathcal{P}_{C}\times\{0,1\}$ to $L$) to a map landing inside the tubular neighbourhood $U$ of $L$. Such a homotopy exists by homotoping $\gamma$ (relative to its endpoints) through $Z_{C}$ to the straight-line path from $x_{0}$ to $x_{0}+i$, and using a map defined with the same formula (with respect to this family of paths instead of $\gamma$) as $T$. Note that here we apply Proposition 4.1. ∎
Proof of Theorem 1.7.
Follows from Lemmas 6.1 and 6.3. ∎